How much water should be added to 8ml of 6% saline solution to reduce the concentration to 4% saline
Answer by richwmiller(17219)

8*.94+x=.96(8+x)
1*x+0.94*8=0.96(8+x)
1*x+7.52=7.68+0.96x
1x-0.96x=7.68-7.52
0.04x=0.16
x=4 ml of water
check
1*4+0.94*8=0.96(8+4)
4+7.52=0.96(12)
11.52=11.52
ok
8*.06=.04(8+x)
.48=.32+.04x
.16=.04x
x=4 ml same answer